# GlucOS: Security, correctness, and simplicity for automated insulin delivery

### v1

@@ 8 Formal verification
A single error or vulnerability in an automated insulin delivery system can have life-threatening consequences for the user. This criticality motivates our use of formal verification in our development and testing process. Through formal verification, we aim to confirm two critical aspects of our insulin delivery system. First, we validate the goals and claims behind our critical insulin delivery functions. Second, we verify that GlucOS applies the correct state transitions in precarious scenarios to protect the user when the system loses confidence in one or more of its components due to miscalculation or malicious behavior.
Systematically identifying the invariants and postconditions in alignment with our claims that we verify for a critical function in our system is important. Our first step is to clearly define what the function is supposed to do. This helps list all the inputs that the function takes, including their types and ranges. Then, we define the expected outputs, including their types and valid ranges. Determining safety and domain-specific constraints for each function plays a major role in ensuring that the function never produces an unexpected output that puts a user at risk. For example, a general constraint would be that a basal rate should never be negative as it is invalid. A domain-specific constraint is defining a maximum basal rate threshold that is in accordance with medical guidelines. For a full list of the invariants that we define and prove, see Appendix D.
These steps help define the preconditions, postconditions and invariants that help prove the properties of our verified functions. We run formal verification in Dafny to verify our claims on how a function works. After we verify a function, we port the code to Swift to make it part of the BioKernel.
We verify invariants and postconditions for the system’s critical functions in the aforementioned approach. For instance, we verify the microbolusing function to ensure a minimum time interval between consecutive micro-boluses to prevent frequent dosing in short intervals (within 4 minutes), confirm that the user glucose level is significantly above the target(atleast 20 mg/dL) before administering a bolus and that the bolus amount is within safe and predefined limits. These are claims about our function that we formally verify to assure that they are always correct.
In another example, we prove correct our insulin delivery calculations to ensure that we maintain accurate accounting of insulin delivery. For this verification task, we verify our insulin delivery over a duration of time function to guarantee non-negative results, correctly handle zero duration or non-overlapping time segments, making sure that the dosage does not exceed recommended limits, and accurately calculate insulin delivery based on time segment overlap.
We perform this type of verification on all the critical functions involved in the insulin delivery process. This high-level of assurance helps us trust the the decisions that GlucOS makes.
We also formally verify the state machine we use for defining system transitions under undesirable scenarios (Section 7). We discover several potential issues in the state transition logic and function design during verification. For example, there were cases where the system could transition to an unsafe state under certain conditions, potentially leading to excessive insulin delivery. We also identified bugs related to violating valid bounds and domain-specific constraints. A more detailed discussion of the bugs have been included as part of subsection  10.6 in our evaluation.
9 Implementation
In our implementation, the BioKernel is an iOS app on iPhone or iPad hardware. It communicates with a Dexcom G7 CGM and Omnipod DASH insulin pump via Bluetooth low energy and runs the core closed-loop algorithm. For the BioKernel, we use the Swift programming language, the Actor abstraction for data-race-free concurrency, and JSON files stored on disk to capture persistent state.
Although the BioKernel is a fully featured and stand alone automated insulin delivery system, GlucOS uses an event logging system that enables separate apps to implement complementary features while remaining isolated from the BioKernel. We implement our event logging system using a Google Cloud App Engine app and the BioKernel produces logs that other apps consume. The other apps include an app for alerting when the individual is likely to experience hypoglycemia or hyperglycemia, an app to replay closed-loop runs to verify the results, and an app for running simple ML on data to calculate therapeutic settings.
To keep predictive ML and the pump driver isolated, we use the Swift type system and object boundaries to maintain isolation. Our architecture also supports the use of process boundaries for even strong isolation and to implement OS-level sandboxing. iOS lacks support for proper process APIs (fork, exec etc), so we use the Swift type system instead.
For CGM and insulin pump drivers, we use LoopKit [41] and isolate untrusted portions using the Swift type system.
@@ 10 Evaluation
Given the risks associated with running ML for automated insulin delivery, we answer most questions on 21 virtual humans from an FDA-approved simulator [9]. The simulator generates CGM readings across different age groups including children, adolescents and adults. We also evaluate security mechanisms on virtual patients to see if they can generalize to individuals with different physiologies. We present experiences of running GlucOS with a real individual in the next section.
10.1 Does ML lead to better outcomes for managing T1D?
We confirm findings from prior research that ML leads to better outcomes [47, 66, 34] by comparing the average proportion of time spent by virtual patients in range when using our reactive safe model (Section 4.2) against our predictive ML model (Section 4.3). As mentioned in Section 2, the American Diabetes Association (ADA) recommends that an individual with T1D should spend more than 70% of a day (also called TIR) [8, 21] in order to remain healthy.
We train separate predictive ML models (feed-forward neural networks trained to minimize MSE loss) for each virtual patient and similarly tune the physiological parameters of the reactive safe model to suit each patient. We ensure that we run the reactive safe model and the predictive ML model on identical conditions within the simulator.
Patient TIR with TIR with
type ML model reactive safe model
Adults 93.49% 88.25%
Adolescents 85.23% 81.98%
Children 86.6% 77.31%
@@ 10.1 Does ML lead to better outcomes for managing T1D?
Figure 8: Average proportion of time spent in range by 21 virtual patients in identical scenarios with a our predictive ML model and our reactive safe model.
We show the average proportion of time spent in range by virtual patients from each age group (adults, adolescents and children) in Figure 8. We can see that ML helps virtual patients across all age groups spend a larger amount of time in range, thereby showing that ML leads to better outcomes for managing T1D.
10.2 What are the risks associated with untrusted ML algorithms for managing T1D?
@@
We assess the impact of predictions made by intentionally malicious ML algorithms on managing T1D. Since GlucOS is designed to be flexible with any predictive algorithm, it could be subject to attacks from malicious ML algorithms. We intentionally introduce an order of magnitude difference to the amount of insulin predicted by the predictive ML model in Section 10.1. In one experiment, we dose one-tenth of the amount of insulin and in another experiment, we dose ten times the amount of insulin. We run these malicious ML algorithms within the simulator on each virtual patient on the same scenarios as the previous experiment.
@@
Figure 9 and Figure 10 show the CGM readings on 3 randomly chosen virtual patients on a particular scenario for the experiments where we dose less and dose more respectively. When dosing less, we see that all patients spend most of their time above Level 1 hyperglycemia which would have long-term effects on their well-being. When dosing more, all patients fully go below Level 2 hypoglycemia, which would have killed them.
@@
Figure 9: Variation in CGM readings for 3 virtual patients when running a predictive ML algorithm that intentionally doses one-tenth the amount of the required insulin. The CGM readings staying mostly above the Level 1 hyperglycemic range would have drastic effects on well-being.
@@
Figure 10: Variation in CGM readings for 3 virtual patients when running a predictive ML algorithm that intentionally doses ten times the amount of the required insulin. The CGM readings staying below the Level 2 hypoglycemic range would have likely been fatal to all 3 patients.
@@
Across all virtual patients, we see an average drop in TIR from 88.44% with correct dosing to 23.32% when dosing less and an average drop in TIR from 88.44% to 0% when dosing more. Correspondingly, when dosing less, we observe an average increase in the timein hyperglycemia from 3.77% with correct dosing to 65.05%. Similarly, when dosing more, we observe an average increase in the time in hypoglycemia from 0% to 100%.
@@
10.3 What is the impact of GlucOS’s algorithmic security mechanism on managing T1D?
GlucOS’s algorithmic security mechanisms should keep the blood glucose of individuals in range even in presence of malicious ML algorithms. The mechanisms should also ensure that they do not significantly dampen the level of control provided by an efficient algorithm.
Impact on malicious algorithms. We repeat the experiments from the previous section with the same scenarios. We run the same malicious algorithms, but also run the reactive safe model to set bounds on the amount of insulin the malicious algorithms can inject. Figure 11 and Figure 12 show CGM readings on the same 3 random virtual patients on the same scenarios from the experiment in the previous section. When running either malicious model, we see that all patients spend over 70% of their time in range. These results show that GlucOS protects individuals even in presence of intentionally malicious algorithms. On average, across all virtual patients, we still see a drop in TIR from 88.44% with correct dosing to 73.89% when employing safety while dosing less and a drop in TIR from 88.44% to 78.22% when dosing more. While the times spent in range is less than correct dosing, the impact of incorrect dosing is significantly dampened with patients spending the recommended amount of time in range.
Figure 11: Variation in CGM readings for 3 virtual patients when clamping an algorithm that intentionally tries to dose one-tenth the amount of the required insulin. On average, the CGM readings stay in range over 70% of the time.
Figure 12: Variation in CGM readings for 3 virtual patients when clamping an algorithm that intentionally tries to dose ten times the amount of the required insulin. On average, CGM readings stay in range over 70% of the time.
Patient TIR with clamped
type Predictive ML model
Adults 92.88%
Adolescents 82.38%
Children 85.42%
Figure 13: Average proportion of time spent in range by 21 virtual patients in the same scenarios in Section 10.1 when clamping the predictive ML with the reactive safe model.
Impact on effective benign algorithms. We evaluate the impact of GlucOS’s algorithmic security mechanisms on benign algorithms such as the predictive ML model used in Section 10.1. We repeat the experiment of running the predictive ML model on all virtual patients in the same random scenarios while maintaining the same parameters for the clamps and reactive model used for the malicious algorithms evaluation.
Comparing Figure 13 against Figure 8 we see that there is only a slight drop in the average proportion of time spent in range by virtual patients across all age groups when employing GlucOS’s algorithmic security mechanisms. The average proportion of time spent in range being higher than the reactive safe model shows that GlucOS’s security does not come at the cost of the benefits from ML.
@@ 10.3 What is the impact of GlucOS’s algorithmic security mechanism on managing T1D?
10.4 Does GlucOS ensure pump operation while overcoming buggy/malicious pump drivers?
We first evaluate GlucOS’s driver security mechanism in terms of its impact on benign pump drivers that accurately relay pump communications. We then evaluate the impact on detecting additional, modified or dropped communication from a malicious pump driver. Our mechanism should ensure that all communications go through with a benign pump driver and should catch all manipulated communication with a malicious pump driver. We do not use the simulator for this evaluation since it does not incorporate a pump driver.
@@ 10.4 Does GlucOS ensure pump operation while overcoming buggy/malicious pump drivers?
Patient TIR without TIR with
type biological invariant biological invariant
Adults 49.93% 88.92%
Adolescents 55.83% 75.26%
Children 67.41% 77.45%
Figure 15: Average proportion of time spent in range by 21 virtual patients when fluctuating insulin sensitivities. We compare cases with and without our end-to-end security mechanism using the biological invariant.
10.5 Is the biological invariant effective?
We run experiments on the simulator to evaluate if GlucOS’s use of the biological invariant to transition to manual insulin dosing leads to better T1D management. Since the simulator does not provide support to fluctuate insulin sensitivity, we intentionally scale up or scale down values for insulin sensitivity when calculating insulin doses in closed loop. We trigger cases where we scale down the insulin sensitivity to simulate scenarios where the system assumes a lower sensitivity value than the user’s actual sensitivity more frequently since this are more relevant to evaluating the biological invariant.
Across all virtual patients, we record an average TIR of 57.56% when fluctuating insulin sensitivity without incorporating the biological invariant. In contrast, we record an average TIR of 80.54% when incorporating the biological invariant to transition to manual dosing. Figure 15 shows the average TIR across age groups. We clarify that we use the correct insulin sensitivity values to compute bolus doses when dosing manually, since users would know their true insulin sensitivity when taking in manual doses.
10.6 Did formal verification help with correctness?
@@ 11 Using GlucOS in the real world
In this section, we report on our experiences working with one individual, Bob, since November 2023. We show our results for using GlucOS with our reactive safe model closed-loop algorithm as well as with our predictive ML model closed-loop algorithm paired with our security mechanism. We also measure system performance, report on errors, and quantify our implementation simplification efforts.
Overall, our study is a correlation study, where we evaluate one individual using our system and compare segments of time with each other.
@@ 11.1 Ethical considerations
In this section, we outline the steps we take to ensure that we uphold high ethical standards and ensure that the human using our system, who we will call Bob, is safe.
First, one of the co-authors of this paper is a board-certified Endocrinologist, who specializes in T1D. They are both a clinician with an active practice seeing patients living with T1D and a researcher who focuses on research on T1D. They designed the safety protocol that we detail in Appendix A, which defines the testing we conduct to ensure that the individual using our system is safe, and the specific criteria we use to stop the study, if needed.
Second, we submitted our study to our university’s IRB, which determined that our “self evaluation” is exempt from IRB (case number omitted, but available).
Third, Bob increased the frequency of his visits to his Endocrinologist, which includes lab work to confirm that he remained in good overall metabolic health.
Overall, GlucOS system had a large positive impact on the individual who used it. Bob’s lab-based A1C test (a blood test that measures average glucose levels over time), which includes his time using GlucOS, was 5.8%. That is nearly non-diabetic levels of control – healthy individuals will have an A1C of 5.6% or lower. This result is a profound outcome for Bob, and GlucOS had a massive positive impact on Bob’s health and quality of life.
@@ 11.2 Is an ML-based closed-loop system safe and effective in practice?
11.3 How does the BioKernel perform?
We look at two aspects of system performance when evaluating our real-world use of GlucOS. First, we measure how long our closed-loop algorithm takes to run and second, we count how often the closed-loop algorithm runs successfully end-to-end. We use Bob’s iPhone 14 to collect all of the data we show here and these numbers are averages across the entire six month period.
On average, the BioKernel spends 4.3s running the closed loop algorithm, where almost half of that time is dominated by CGM and pump communications. Of the components within the closed-loop algorithm, the ML predictions take 2.2s on average and the safety logic takes 19ms (0.019s). Since the system runs in the background, we believe that iOS runs the device with reduced computational capacity to save power. But overall, 4.3s of wall-clock time for every five minutes our closed-loop algorithm runs with only a tiny fraction of that coming from our safety logic represents practical system performance.
For successful closed-loop end-to-end execution, over the evaluation period the closed-loop algorithm runs successfully 96.5% of the time, with 2.1% of the runs failing due to pump communication failures, which is our largest class of failure.
11.4 How simple is our BioKernel?
To evaluate the complexity of our BioKernel, we count the lines of code in our implementation and compare against Loop, another open-source automated insulin delivery system.
Using the “cloc” utility the BioKernel app has 4.0k lines of code compared to 39.8k lines of code in the Loop app, an order of magnitude reduction. We omit the lines of code coming from LoopKit, which both the BioKernel and Loop use, because our pump security measures remove most of it from our trusted computing base. However, LoopKit and the drivers have a substantial amount of code, weighing in at 70.1k lines of code, showing the importance of our security measures.
To explain why we have such a large reduction in source code, we outline the differences between the two systems. First, Loop has several features that we move outside of the BioKernel. These features include a remote interface for insulin dosing, a Watch app, Siri command interfaces, third party libraries, tutorials, complex meal announcements, and physiological simulations. Second, we simplify our implementation of features that are shared between the BioKernel and Loop. These shared features include a simplified local storage implementation, simplified concurrency support, minimal therapeutic settings, and a simplified closed-loop algorithm. As our evaluation shows, we can still provide tight control over T1D with this simplified core system and still support most or all these additional features architecturally.

### v2

@@ 7 Formal verification
To provide additional assurance that our implementaiton is correct, we port our Swift code to Dafny and use Hoare logic. As the system we are verifying is complex, our goal is to locate the pieces and target verification in places where formal methods can have the greatest impact; this is a common reality of applying formal methods at scale, see for example [19]. Since we are porting the Dafny code to Swift, to reduce the chance of human error, we use a line-by-line port as much as possible and add unit tests for each function to provide confidence that the two implementations are equivalent. We port updated code back to Swift manually because Dafny does not include code generation for Swift.
These efforts paid off: we found and fixed nine bugs in our original implementation (Appendix D); in addition to proving various invariants in the code for our insulin delivery path and security mechanisms. However, the most interesting part was where these bugs were: all of our core algorithms were correct, but we found bugs in seemingly less critical parts of the system, like user settings, which had subtle interactions with the core algorithms. For a full list of the invariants that we define and prove, see Appendix E.
@@ 8 Evaluation
8.1 How effective is GlucOS’s algorithmic security mechanism?
We evaluate GlucOS’s algorithmic security mechanism on 21 virtual humans from an FDA-approved simulator [7]. We use simulation since this evaluation involves injecting dangerous amounts of insulin.
To be effective, GlucOS’s algorithmic security mechanism should protect users against untrusted ML algorithms that administer incorrect insulin doses (overdose or underdose), while minimally impeding algorithms that are efficient at managing T1D. Accordingly, we first run simulations with a predictive ML model inspired from prior research [52], without incorporating our security mechanism to set a baseline of an algorithm that is efficient at managing T1D. Then, in the same simulation scenarios, we introduce an order of magnitude difference (increase and decrease) to the insulin doses computed by the baseline algorithm, without incorporating our security mechanism to quantify the risks of untrusted algorithms. We then repeat both sets of simulations while incorporating our security mechanism to quantify the impact of the mechanism on correct and incorrect dosing.
The goal of T1D management is to maintain glucose levels that are in a safe, target range (70-180 mg/dl) for the majority of the time – an outcome that has been linked to positive long-term health outcomes [53]. Glucose levels 54-69 mg/dl (level 1 hypoglycemia) or less than 54 mg/dl (level 2 hypoglycemia) put individuals at risk for acute complications like impaired consciousness, seizure, and death, while levels 181-250 mg/dl (level 1 hyperglycemia) or greater than 250 mg/dl (level 2 hyperglycemia) increase risks for long-term vascular damage. The American Diabetes Association (ADA) recommends that individuals spend at least 70% of their time in range [6, 24] in order to remain healthy.
8.1.1 Without algorithmic security mechanism
Predictive ML model baseline With the predictive ML model, we followed all ML best practices in carefully tuning hyperparameters and evaluating the model on simulation scenarios that were different from those used in training. Figure 7 shows the average proportion of time spent in range (TIR) by virtual humans from different age groups (adults, adolescents and children) when employing the predictive ML model. We see that individuals across all groups spend over 85% of their time in range compared to 77% when using only the reactive safe model. These results provide a baseline for an algorithm that is effective in managing T1D.
Reactive TIR with TIR with
safe predictive clamped
Age group model ML model ML model
Adults 88.25% 93.49% 92.88%
Adolescents 81.98% 85.23% 82.38%
Children 77.31% 86.6% 85.42%
@@ 8.1.1 Without algorithmic security mechanism
Figure 7: Average proportion of time spent in range (TIR) across 21 virtual humans with a reactive safe model, a predictive ML model, and clamping the predictive ML with the reactive safe model for security.
@@
Incorrect ML dosing Since GlucOS is designed to be flexible with any predictive algorithm, it could be subject to attacks from malicious ML algorithms. To simulate attacks, we intentionally introduce an order of magnitude difference to the amount of insulin predicted by the baseline predictive ML model. In one experiment, we dose one-tenth of the amount of insulin and in another experiment, we dose ten times the amount of insulin computed by the predictive ML model. We run these malicious ML algorithms within the simulator on each virtual human on the same scenarios as the previous experiment.
@@
Figure 8 and Figure 9 show the CGM readings on three randomly chosen virtual humans on a particular scenario for the experiments where we dose less and dose more respectively. When dosing less, we see that all of them spend most of their time above Level 1 hyperglycemia. When dosing more, all individuals fully go below Level 2 hypoglycemia, which would have killed them.
@@
Figure 8: Variation in CGM readings for three virtual humans when running a predictive ML algorithm that intentionally doses one-tenth the amount of the required insulin. The CGM readings staying mostly above the Level 1 hyperglycemic range.
@@
Figure 9: Variation in CGM readings for three virtual humans when running a predictive ML algorithm that intentionally doses ten times the amount of the required insulin. The CGM readings staying below the Level 2 hypoglycemic range would have likely been fatal to all three individuals.
@@
Across all virtual humans, we see an average drop in TIR from 88.44% with correct dosing to 23.32% when dosing less and an average drop in TIR from 88.44% to 0% when dosing more. Correspondingly, when dosing less, we observe an average increase in the time in hyperglycemia from 3.77% with correct dosing to 65.05%. Similarly, when dosing more, we observe an average increase in the time in hypoglycemia from 0% to 100%.
@@
8.1.2 With algorithmic security mechanism
Impact on incorrect dosing. We repeat the experiments from the previous section with the same scenarios. We run the same models for incorrect dosing, but also run our reactive safe model (Section 3) to set bounds on the amount of insulin the malicious algorithms can inject.
Figure 10 and Figure 11 show CGM readings on the same three virtual humans on the same scenarios from the experiment in the previous section. With both types of incorrect dosing, we see that all of them spend over 70% of their time in range. These results show that GlucOS protects individuals even in presence of malicious incorrect dosing. On average, across all virtual humans, we only see a drop in TIR from 88.44% with correct dosing to 73.89% when employing safety while dosing less and a drop in TIR from 88.44% to 78.22% when dosing more. While the times spent in range is less than that of correct dosing, the impact of incorrect dosing is significantly dampened with individuals spending the recommended amount of time in range, despite facing an active attack.
Figure 10: Variation in CGM readings for three virtual individuals when clamping an algorithm that intentionally tries to dose one-tenth the amount of the required insulin. On average, the CGM readings stay in range over 70% of the time.
Figure 11: Variation in CGM readings for three virtual humans when clamping an algorithm that intentionally tries to dose ten times the amount of the required insulin. On average, CGM readings stay in range over 70% of the time.
Impact on effective benign algorithms. We evaluate the impact of GlucOS’s algorithmic security mechanism on benign algorithms that are effective in managing T1D such as the baseline predictive ML model. We repeat the experiment of running the predictive ML model on all virtual humans in the same random scenarios while maintaining the same parameters for the clamps and reactive model used to evaluate incorrect dosing.
Comparing data in Figure 7 we see that there is only a slight drop in the average proportion of time spent in range by virtual humans across all age groups when employing GlucOS’s algorithmic security mechanism. When employing the reactive safe model as is to calculate insulin doses, we report an average TIR of 82.48% which is lower than the average TIR of 86.84% observed when running the predictive ML model with GlucOS’s security mechanism. This shows that GlucOS’s algorithmic security guarantees do not come at the cost of effective T1D management from predictive ML.
@@ 8.1.2 With algorithmic security mechanism
8.2 Does GlucOS ensure pump operation while overcoming malicious pump drivers?
We first evaluate GlucOS’s driver security mechanism in terms of its impact on benign pump drivers that accurately relay pump communications. We then evaluate the impact on detecting additional, modified, or dropped communication from a malicious pump driver. Our mechanism should ensure that all communications go through with a benign pump driver and should catch all manipulated communication with a malicious pump driver. We do not use the simulator for this evaluation since it does not incorporate a pump driver.
@@ 8.2 Does GlucOS ensure pump operation while overcoming malicious pump drivers?
Age Group TIR without
biological invariant
Adults 49.93%
Adolescents 55.83%
Children 67.41%
Figure 13: Average proportion of time spent in range across 21 virtual humans when fluctuating their insulin sensitivity. The low values for TIR stress the importance of an end-to-end security mechanism, i.e., the biological invariant.
8.3 Is the biological invariant important?
We run experiments on the simulator to assess the risks of not considering our end-to-end security mechanism, i.e., the biological invariant. Figure 13 shows the average TIR across age groups. Across all virtual humans, we record an average TIR of 57.56%, even with our algorithmic security mechanism, when fluctuating insulin sensitivity. This corresponds to an average drop of 29.28% as a result of not considering the biological invariant. These results emphasize the importance of incorporating an end-to-end invariant and undermines results obtained from pure simulation. We present case-studies involving the biological invariant with a real user using GlucOS in Section 9.3.
@@ 9 Using GlucOS in the real world
In this section, we report on our experiences working with seven individuals using GlucOS to help manage their T1D. Six people are students who only used the predictive alerting feature of the BioKernel (Section 6) for two months, from June 2024 - August 2024. One individual, who we will call Bob, has been using the full GlucOS system since November 2023 for insulin delivery and predictive alerting.
Our user study for the six students was approved by our university’s IRB, and Bob’s use GlucOS was determined by our university’s IRB to be “self evaluation” and exempt from IRB.
In our implementation, the BioKernel is an iOS app that consists of 5.1k lines of code in our trusted computing base. It communicates with a Dexcom G6 or G7 CGM and Omnipod DASH insulin pump via Bluetooth low energy, and runs the core closed-loop algorithm. For the BioKernel, we use the Swift programming language, the Actor abstraction for data-race-free concurrency, and JSON files stored on disk to capture persistent state. Although the BioKernel is a fully featured and stand alone automated insulin delivery system, GlucOS uses an event logging system that enables separate apps to implement complementary features while remaining isolated from the BioKernel.
@@ 9.1 Ethical considerations
In this section, we outline the steps we take to ensure that we uphold high ethical standards and ensure that the people using our system, are safe.
One of the co-authors of this paper is a board-certified Endocrinologist, who specializes in T1D. They are both a clinician with an active practice seeing patients living with T1D and a researcher who focuses on research on T1D. They designed the safety protocol, which defines the testing we conduct to ensure that the individuals using our system are safe, and the specific criteria we use to stop the study, if needed. Appendix A includes a more detailed discussion of our ethical considerations and safety protocol.
@@ 9.2 Is an ML-based closed-loop system safe and effective in practice?
9.3 Do biological invariant violations trigger in practice?
To determine if biological invariant violations trigger in practice, we review 90 days of data for Bob, from May 29th, 2024 to August 27th, 2024. During this period of time, biological invariant violations occurred 1.6k times, or 9.0% of the time. Of these violations there were 49 sequences of 30 minutes or longer with the longest sequence being 65 minutes.
These results suggest that although natural variations in glucose metabolism do lead to biological invariant violations, none of them were long enough to trigger our most severe action of transitioning to manual mode after two hours.
We reviewed one weeks worth of violations with Bob manually to determine the cause. During the most recent week of data, Bob had five sequences of 30 minutes or longer and all were caused by exercise.
9.4 Are predictive alerts useful for security?
To measure predictive alerts as a security mechanism (Section 6), we recruited six students who live with T1D and had them use the GlucOS predictive alerts to complement their default CGM alerts. Our primary metric is time spent below 70 mg/dl, where a decrease in the amount of time spent in hypoglycemia suggests that predictive alerts are effective to empower people to mitigate security attacks.
For the two month study, our participants decreased their average time in hypoglycemia from 2.7% down to 1.6%, and all individuals experienced a decrease in the amount of time spent in hypoglycemia. For context, participants in a state-of-the-art automated insulin delivery clinical trial experienced a decrease of time in hypoglycemia of 0.9% [15], showing that our 1.1% decrease is substantial.
Overall, our results suggest that predictive alerts can be a useful last line of defense for trustworthy automated insulin delivery systems.